OpenStudy (anonymous):

From Y(s)(s^2+2s+2)=alpha/(s^2+(alpha)^2), how can you leave Y(s) by itself? What does it look like?

OpenStudy (anonymous):

just multiply both sides by \[\frac{ 1 }{ s^{2}+2s+2 }\] so you would get : \[Y(s)=\frac{ \alpha }{ (s^{2}+2s+2)(s^{2}+\alpha^{2}) }\] o.O are you seriously asking for this ?

OpenStudy (anonymous):

Then you take the inverse laplace transform and it looks like alpha/(s^2+2s+2) and alpha/(s^2+(alpha)^2), right?

OpenStudy (anonymous):

um i got no idea what u did there , and got no idea what ur asking xD

OpenStudy (anonymous):

Like you supposed to find the inverse laplace transform of alpha/((s^2+2s+2)(s^2+(alpha)^2), but how would you simplify this?

OpenStudy (anonymous):

um i dont think u need to simplify anything because it looks fine like that i am just not sure how to do it because of the alpha in it .. .

OpenStudy (anonymous):

Okay, got it.

OpenStudy (loser66):

hey, stubborn girl. What is the original one?

OpenStudy (loser66):

I mean the form of y" +....